# Hyper-Transforming Latent Diffusion Models

### v2

@@ 3.2.1 Hyper-Transformer Decoder
The Hyper-Transformer Decoder (HD) is depicted in Figure 3. Mathematically, it is defined as a function of the latent tensor that produces the parameters $\Phi=g_{\phi}({\bm{z}})$ of the INR. We elaborate on previous work that proposed Vision Transformers Dosovitskiy (2020) for meta-learning of INR parameters from images Chen and Wang (2022), to design an efficient full Transformer architecture Vaswani (2017) that processes latent variables into INR parameters.
@@
Tokenizer The HD decoder begins by splitting tensor-shaped latent variables into $N$ patches ${\bm{z}}_{p}\in\mathbb{R}^{P^{2}\times d_{z}}$ of fixed size $(P,P)$, where $N=H_{z}\cdot W_{z}/P^{2}$. Each patch is then flattened and projected into a lower-dimensional embedding space using a shared linear transformation. The embedding for each patch serves as an input token for the Transformer.
@@
Transformer Encoder Following tokenization, the first half of the HD decoder, a Transformer Encoder, applies multi-head self-attention mechanism, repeated for several layers, outputting tokenized embeddings that we referred to as latent tokens.
@@
Transformer Decoder Unlike prior work Chen and Wang (2022) that only employs Transformer Encoders, we introduce a Transformer Decoder that cross-attends to latent tokens to generate output tokens, which are then mapped to the column weights of the INR weight matrices. The input to the Transformer Decoder consists of a globally shared, learnable set of initial weight tokens, denoted as $\bar{{\mathbf{W}}}^{\text{i}}$, which serve as queries. The decoder processes these queries via cross-attention with the latent tokens, which act as keys and values. The final set of column weights, ${\mathbf{W}}^{\text{o}}$, is obtained by applying a linear transformation to the output tokens, ensuring they match the corresponding column dimensions of the INR weight matrices. The biases of the INR are modelled as globally shared, learnable parameters, as learning the weight space alone provides sufficient flexibility to model the stochastic process while simplifying the task of the HD decoder.
@@
Weight Grouping Following Chen et al. (2021), we adopt a similar weight grouping strategy to balance precision and computational efficiency in INR parameter generation. However, our approach differs in the reconstruction strategy applied to the Transformer-generated weights.
@@
Let ${\mathbf{W}}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ represent a weight matrix of an INR layer (omitting the layer index $l$ for simplicity), where each column is denoted as ${\bm{w}}_{c}$ for $c\in\{1,\dots,d_{\text{in}}\}$. Directly mapping a Transformer token to every column is computationally expensive; instead, we define $G$ groups per weight matrix, where each group represents $k=\frac{d_{\text{in}}}{G}$ columns (assuming divisibility). This results in ${\mathbf{W}}^{\text{o}},\bar{{\mathbf{W}}}^{\text{i}}\in\mathbb{R}^{d_{\text{out}}\times G}$, since ${\mathbf{W}}^{\text{o}}$ and $\bar{{\mathbf{W}}}^{\text{i}}$ have the same dimensions. The full weight matrix is then reconstructed as:
@@
where $\mathcal{R}$ denotes the reconstruction operator, ${\bm{w}}_{c}$ is the $c$th column of ${\mathbf{W}}$, and $\bar{{\bm{w}}}_{c}^{\text{b}}$ is the corresponding column of $\bar{{\mathbf{W}}}^{\text{b}}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$, a set of learnable base parameters that provide an initial weight structure. In Trans-INR Chen and Wang (2022), the Transformer-based hypernetwork is limited to generating normalized weights using the following reconstruction method:
@@
However, we found this approach unsuitable for generating weights in INRs with periodic activations. We hypothesize that the normalization in Eq. (8) causes the resulting weights (and their gradients) to vanish when $||{\bm{w}}_{\lfloor c/k\rfloor}^{\text{o}}||\approx 0$, leading to training instability.
To address this, we propose a new reconstruction operator that removes the constraint on weight normalization, yielding significantly more stable training, particularly for INRs with periodic activations:
@@
This formulation ensures that when $||{\bm{w}}_{\lfloor c/k\rfloor}^{\text{o}}||\approx 0$, the base weights remain unchanged, preventing unwanted instability in the generated INR parameters.
It is essential to distinguish between the two sets of globally shared learnable parameters. The sequence $\bar{{\mathbf{W}}}^{\text{i}}$ (of length $G$) serves as the initialization for the grouped weight tokens and is provided as input to the Transformer Decoder. In contrast, the sequence $\bar{{\mathbf{W}}}^{\text{b}}$ (of length $d_{in}$) acts as a global reference for the full set of INR weights and is used exclusively for the reconstruction in Eq. (9).
@@ 3.3 Variational Inference
Given our implicit modelling of INR parameters, the objective is to learn a structured and compact probabilistic latent space $p_{\theta}({\bm{z}})$ that enables meaningful posterior inference via Bayesian principles. Specifically, we seek to approximate the true posterior $p({\bm{z}}|{\bm{X}},{\bm{Y}})$, facilitating accurate and flexible inference.
To achieve this, we extend the Variational Autoencoder (VAE) framework Kingma (2013) to the efficient generation of INRs, which we refer to as I-VAE. Further, we incorporate Latent Diffusion Models to enhance the expressiveness of the latent space. Our full model is depicted in Figure 2, and the following sections describe its key components.
@@
During the first stage of training, following LDMs Rombach et al. (2022), we impose a simplistic standard Gaussian prior $p({\bm{z}})=\mathcal{N}(\bm{0},\bm{I})$, and we set a low $\beta$ value to encourage high reconstruction accuracy while promoting a structured latent space that preserves local continuity. This choice facilitates smooth interpolations and improves the quality of inferred representations.
In addition, following Rombach et al. (2022), for the case of images, we incorporate perception losses Zhang et al. (2018), and patch-based Isola et al. (2017) adversarial objectives Dosovitskiy and Brox (2016); Esser et al. (2021); Yu et al. (2022).
@@ 3.4 Latent Diffusion
In this stage, we leverage the pre-trained I-VAE and the highly structured latent space encoding INRs to fit a DDPM Ho et al. (2020) to the aggregate posterior:
@@
where $p_{\theta}({\bm{z}})$ represents the learned diffusion-based prior over the latent space. As shown by Ho et al. (2020), the variational lower bound can be further transformed into a Denoising Score Matching (DSM) objective Vincent (2011)
@@
3.5 Hyper-Transforming Latent Diffusion Models
As an alternative to full training, we introduce a highly efficient approach for training LDMI when a pre-trained LDM for discretized data is available. In this setting, we eliminate the need for two-stage training by leveraging the structured latent space learned by the pre-trained VAE. Specifically, we freeze the VAE encoder and the latent-diffusion model and train only the Hyper-Transformer Decoder to maximize the likelihood of the decoded outputs, using the objective
$\mathcal{L}_{\text{HT}}(\phi)=\mathbb{E}_{q_{\psi}({\bm{z}}|{\bm{X}}_{m},{\bm{Y}}_{m})}\left[\log p_{\Phi}({\bm{Y}}|{\bm{X}})\right],$ (15)
to which we can add perceptual and adversarial objectives. The HD decoder is flexible enough to extract the information from the pre-learned latent space.
We refer to this training strategy as hyper-transforming, as it efficiently adapts a standard LDM to an INR-based framework without requiring retraining of the latent encoder or prior. The full procedure is detailed in Algorithm 1.
4 Experiments
4.1 Training Setup
Depending on the nature of the data, we utilize different architectures for the encoder. For image and climate data, we employ ResNets He et al. (2016), while for point clouds, we employ 3D-convolutional networks.
We evaluate our proposed model on datasets spanning multiple domains: (1) natural image datasets, including CelebA Liu et al. (2015) and ImageNet Russakovsky et al. (2015); (2) 3D objects, specifically the Chairs subclass from the ShapeNet repository Chang et al. (2015), which provides approximately 6,778 chair models for 3D reconstruction and shape analysis; and (3) polar climate data, using the ERA5 temperature dataset Hersbach et al. (2019), for analyzing global climate dynamics. In our hyper-transforming setup, we leverage publicly available pre-trained Latent Diffusion Models from Rombach et al. (2022), specifically the LDM-VQ-4 variant trained on CelebA-HQ at $(64\times 64)$ resolution, and the LDM-VQ-8 variant trained on ImageNet.
@@ 4.1 Training Setup
(b) ImageNet
Figure 4: (a) Uncurated samples from GASP, Functa (diffusion-based) trained with CelebA-HQ at 64$\times$64 resolution, and our model LDMI trained on CelebA (64$\times$64). (b) Uncurated class-conditional samples from our model LDMI trained on ImageNet (256$\times$256) for a few epochs using the hyper-transforming approach.
@@
Figure 5: Reconstructions from Functa (diffusion-based), VAMoH, and our LDMI compared to ground truth (GT) across four datasets (a)–(d). In (a), the baselines were trained on CelebA-HQ, a zoomed-in 64-sized version of CelebA with more details. Interestingly, our LDMI, trained on the original zoomed-out images at 64$\times$64, produces finer details. In (d), LDMI was trained by hyper-transforming for a few epochs.
@@ 4.2 Generation
GASP (Dupont et al., 2022a) - 7.42 25.7M
Functa (Dupont et al., 2022b) $\leq$ 30.7 40.40 -
@@
Model Chairs (PSNR) $\uparrow$ ERA5 (PSNR) $\uparrow$
Functa (Dupont et al., 2022b) 29.2 34.9
@@
In Figure 4, we provide samples from CelebA at 64$\times$64 and class-conditional samples from ImageNet at 256$\times$256, demonstrating our models ability to generate high-quality and diverse images across different scales, and the qualitative superiority against the baselines. Additionally, in Table 1, we report FID scores of our samples for CelebA and ImageNet, highlighting the model’s performance in terms of image quality and reconstruction accuracy. Importantly, while GASP achieves a lower FID score on CelebA-HQ, it cannot perform reconstructions due to its adversarial design. In contrast, our LDMI supports both high-quality sampling and accurate reconstructions, as demonstrated in the following experiment.
4.3 Reconstruction
Similarly, our model enables seamless reconstruction at arbitrary resolutions. Figures LABEL:fig:super_reconstructions_celeba_256 and LABEL:fig:super_reconstructions_celeba_64 show results from LDMI trained on CelebA-HQ at $(256\times 256)$ and $(64\times 64)$, respectively. To achieve this, we encode the observations, sample from the posterior distribution, and transform these latent samples into INRs using our HD decoder. The resulting INRs are then evaluated on grids of varying resolutions to produce reconstructions ranging from ×$\times 1/8$ to ×$\times 4$. Importantly, such flexible resolution control is not supported by methods like GASP, which cannot perform reconstructions due to their adversarial design.
Figure 5 presents a qualitative comparison between original images from the test split and their reconstructions by our model. As shown in the figure, our model successfully preserves fine details and global structures across different images. For CelebA-HQ $64\times 64$, we achieve a PSNR value of $24.80$ dB, which compares competitively to the previous methods. The high PSNR score suggests there is minimal information loss during the encoding-decoding process while maintaining perceptual quality. On ImageNet, our model demonstrates robust reconstruction performance across diverse object categories and scenes, as shown in Figure 5d, including textures, object boundaries and colour distributions.
Quantitative evaluations in Table 1 indicate competitive performance compared to existing INR-based generative models. It is important to note several key distinctions when interpreting the comparative results in this Table. While Functa exhibits higher PSNR values, this advantage stems from its test-time optimization procedure—fitting a separate modulation vector per test image using ground truth information—rather than the amortized inference approach employed by our model. This fundamental methodological difference undermines a direct comparison, though we include these results for completeness.
@@ 4.4 Hyper-Transforming
On ImageNet and CelebA-HQ at $256\times 256$, our LDMI was trained leveraging the hyper-transforming approach. Notably, despite being trained for only a limited number of iterations compared to other alternatives, the results reported in Figures 4b and 5d already exhibit promising structure and diversity. Given that prior works on INR-based generation require extensive training schedules to achieve high-quality synthesis, we expect significant improvements in fidelity and coherence with extended training.
Our approach benefits from the structured latent space learned by the pre-trained LDM, allowing for efficient adaptation while preserving global consistency. These findings suggest that hyper-transforming provides a scalable and effective strategy for INR-based synthesis, particularly for large-scale datasets like ImageNet. The dual capability of high-quality generation and accurate reconstruction positions LDMI as a versatile foundation for various downstream tasks requiring both generative and reconstructive abilities, a unique advantage over specialized alternatives optimized for only one of these objectives.

### v3

@@ 3.2.1 Hyper-Transformer Decoder
The Hyper-Transformer Decoder (HD) is a central component of our autoencoding framework for generating INRs, referred to as I-VAE and illustrated in Figure 2. Its internal architecture is detailed in Figure 3. Mathematically, it is defined as a function of the latent tensor that produces the parameters $\Phi=g_{\phi}({\bm{z}})$ of the INR. We elaborate on previous work that proposed Vision Transformers (Dosovitskiy et al., 2021) for meta-learning of INR parameters from images (Chen & Wang, 2022), to design an efficient full Transformer architecture (Vaswani et al., 2017) that processes latent variables into INR parameters.
Tokenizer
@@
The HD decoder begins by splitting tensor-shaped latent variables into $N$ patches ${\bm{z}}_{p}\in\mathbb{R}^{P^{2}\times d_{z}}$ of fixed size $(P,P)$, where $N=H_{z}\cdot W_{z}/P^{2}$. Each patch is then flattened and projected into a lower-dimensional embedding space using a shared linear transformation. The embedding for each patch serves as an input token for the Transformer.
Transformer Encoder
@@
Following tokenization, the first half of the HD decoder, a Transformer Encoder, applies multi-head self-attention mechanism, repeated for several layers, outputting tokenized embeddings that we referred to as latent tokens.
Transformer Decoder
@@
Unlike prior work (Chen & Wang, 2022) that only employs Transformer Encoders, we introduce a Transformer Decoder that cross-attends to latent tokens to generate output tokens, which are then mapped to the column weights of the INR weight matrices. The input to the Transformer Decoder consists of a globally shared, learnable set of initial weight tokens, denoted as $\bar{{\mathbf{W}}}^{\text{i}}$, which serve as queries. The decoder processes these queries via cross-attention with the latent tokens, which act as keys and values to obtain the output set of column weights, ${\mathbf{W}}^{\text{o}}$.
Weight Grouping
@@
Following Chen et al. (2021), we adopt a similar weight grouping strategy to balance precision and computational efficiency in INR parameter generation. However, our approach differs in the reconstruction strategy applied to the Transformer-generated weights.
@@
Let ${\mathbf{W}}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ represent a weight matrix of an INR layer (omitting the layer index $l$ for simplicity), where each column is denoted as ${\bm{w}}_{c}$ for $c\in\{1,\dots,d_{\text{in}}\}$. Directly mapping a Transformer token to every column is computationally expensive, as it would require processing long sequences. Instead, we define $G$ groups per weight matrix, where each group represents $k=\frac{d_{\text{in}}}{G}$ columns (assuming divisibility). This results in a Transformer working with grouped weight matrices $\bar{{\mathbf{W}}}^{\text{i}},{\mathbf{W}}^{\text{o}}\in\mathbb{R}^{d_{\text{out}}\times G}$, since $\bar{{\mathbf{W}}}^{\text{i}}$ (input tokens) and ${\mathbf{W}}^{\text{o}}$ (output tokens) have the same dimensions. The full weight matrix is reconstructed as:
@@
where $\mathcal{R}$ is the reconstruction operator, ${\bm{w}}_{c}$ denotes the $c$th column of ${\mathbf{W}}$, and $\bar{{\bm{w}}}_{c}^{\text{b}}$ is the corresponding column of $\bar{{\mathbf{W}}}^{\text{b}}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$—a set of learnable base parameters that serve as a weight template.
In Trans-INR (Chen & Wang, 2022), the Transformer-based hypernetwork generates normalized weights using the following reconstruction method:
@@
However, we found this approach unsuitable for generating weights in INRs with periodic activations. We hypothesize that the normalization in Equation (8) causes the resulting weights and their gradients to vanish when $||{\bm{w}}_{\lfloor c/k\rfloor}^{\text{o}}||\approx 0$, leading to training instability.
To address this, we introduce a new reconstruction operator that removes the normalization constraint, yielding significantly more stable training, particularly for INRs with periodic activations:
@@
This formulation ensures that when $||{\bm{w}}_{\lfloor c/k\rfloor}^{\text{o}}||\approx 0$, the base weights remain unchanged, preventing instability in the generated INR parameters.
It is essential to distinguish between the two sets of globally shared learnable parameters. The sequence $\bar{{\mathbf{W}}}_{1:L}^{\text{i}}$ (of length $G\times L$) serves as the initialization for the grouped weight tokens and is provided as input to the Transformer Decoder. In contrast, the sequence $\bar{{\mathbf{W}}}_{1:L}^{\text{b}}$ (of length $d_{in}\times L$) acts as a global reference for the full set of INR weights and is used exclusively for the reconstruction in Equation (9).
@@ 3.3 Variational Inference
Given our implicit modelling of INR parameters, the objective is to learn a structured and compact probabilistic latent space $p_{\theta}({\bm{z}})$ that enables meaningful posterior inference via Bayesian principles. Specifically, we seek to approximate the true posterior $p({\bm{z}}|{\bm{X}},{\bm{Y}})$, facilitating accurate and flexible inference. To achieve this, we extend the Variational Autoencoder (VAE) framework (Kingma & Welling, 2013) to the efficient generation of INRs, which we refer to as I-VAE (depicted in Figure 2).
@@
During the first stage of training, following LDMs (Rombach et al., 2022), we impose a simplistic standard Gaussian prior $p({\bm{z}})=\mathcal{N}(\bm{0},\bm{I})$, and we set a low $\beta$ value to encourage high reconstruction accuracy while promoting a structured latent space that preserves local continuity. This choice facilitates smooth interpolations and improves the quality of inferred representations.
In addition, following Rombach et al. (2022), for the case of images, we incorporate perception losses (Zhang et al., 2018), and patch-based (Isola et al., 2017) adversarial objectives (Dosovitskiy & Brox, 2016; Esser et al., 2021; Yu et al., 2022).
@@ 3.4 Latent Diffusion
Figure 4: Schematic of the DDIM-based latent diffusion.
In this stage, we leverage the pre-trained I-VAE and the highly structured latent space encoding INRs to fit a DDPM (Ho et al., 2020) to the aggregate posterior:
@@
where $p_{\theta}({\bm{z}})$ represents the learned diffusion-based prior over the latent space. We instantiate the following variation of the Denoising Score Matching (DSM; Vincent, 2011) objective presented in Equation (1):
@@
(b) CelebA-HQ $256\times 256$
Figure 5: Uncurated samples from GASP, Functa (diffusion-based) and our LDMI trained with CelebA-HQ at 64$\times$64 resolution (a). In (b), our model was trained on CelebA-HQ at 256$\times$256 using the hyper-transforming approach.
3.5 Hyper-Transforming Latent Diffusion Models
As an alternative to full training, we introduce a highly efficient approach for training LDMI when a pre-trained LDM for discretized data is available. In this setting, we eliminate the need for two-stage training by leveraging the structured latent space learned by the pre-trained LDM. Specifically, we freeze the VAE encoder and the LDM and train only the HD decoder to maximize the log-likelihood of the decoded outputs, using the objective
$\mathcal{L}_{\text{HT}}(\phi)=\mathbb{E}_{q_{\psi}({\bm{z}}|{\bm{X}},{\bm{Y}})}\left[\log p_{\Phi}({\bm{Y}}|{\bm{X}})\right],$ (15)
optionally combined with perceptual and adversarial objectives. The HD decoder is sufficiently expressive to extract meaningful information from the pre-trained latent space.
We refer to this strategy as hyper-transforming, as it enables efficient adaptation of a standard LDM to an INR-based framework without retraining the latent encoder or diffusion prior. Notably, this approach is compatible with LDMs based on VAEs, VQVAEs, or VQGANs—since all that is required is access to a pre-trained encoder and diffusion model. The full procedure is outlined in Algorithm 1.
4 Experiments
4.1 Training Setup
Depending on the nature of the data, we utilize different architectures for the encoder. For image and climate data, we employ ResNets (He et al., 2016), while for 3D occupancy data, we employ 3D-convolutional networks.
@@ 4.1 Training Setup
Figure 6: Reconstructions from Functa (diffusion-based), VAMoH, and our LDMI compared to ground truth (GT) across three datasets (a)–(c). In (d) and (e) LDMI was trained by hyper-transforming pre-trained LDMs.
Our evaluations [^1] The code for reproducing our experiments can be found at https://github.com/ipeis/LDMI. span multiple domains: (1) natural image datasets, including CelebA-HQ (Liu et al., 2015) at several resolutions and ImageNet (Russakovsky et al., 2015); (2) 3D objects, specifically the Chairs subclass from the ShapeNet repository (Chang et al., 2015), which provides approximately 6,778 chair models for 3D reconstruction and shape analysis; and (3) polar climate data, using the ERA5 temperature dataset (Hersbach et al., 2019) to analyze global climate dynamics. In our hyper-transforming setup, we leverage publicly available pre-trained Latent Diffusion Models from Rombach et al. (2022), specifically the LDM-VQ-4 variants trained on CelebA-HQ at $256\times 256$ resolution and ImageNet, respectively.
@@ 4.2 Generation
In Figure 5, we provide samples from LDMI trained on CelebA at 64$\times$64 and 256$\times$256, demonstrating its ability to generate high-quality and diverse images across different scales, and the qualitative superiority against the baselines. Additionally, in Table 1, we report FID scores of our samples for CelebA and ImageNet, highlighting the model’s performance in terms of image quality and reconstruction accuracy. Importantly, while GASP achieves a lower FID score on CelebA-HQ, it cannot perform reconstructions due to its adversarial design. In contrast, our LDMI supports both high-quality sampling and accurate reconstructions, as demonstrated in the following experiment.
@@
GASP (Dupont et al., 2022b) - 7.42 25.7M
Functa (Dupont et al., 2022a) $\leq$ 30.7 40.40 -
@@
4.3 Reconstruction
Figure 6 presents a qualitative comparison between original images from the test split and their reconstructions by our model. As shown in the figure, our model successfully preserves fine details and global structures across different images. On ImageNet, our model demonstrates robust reconstruction performance across diverse object categories and scenes, including textures, object boundaries and colour distributions.
Quantitative evaluations in Table 1 indicate competitive performance compared to existing INR-based generative models. For CelebA-HQ $64\times 64$, we achieve a PSNR value of $24.80$ dB, which compares competitively to the previous methods. The high PSNR score suggests there is minimal information loss during the encoding-decoding process while maintaining perceptual quality. It is important to note several key distinctions when interpreting the comparative results in this Table. While Functa exhibits higher PSNR values, this advantage stems from its test-time optimization procedure—fitting a separate modulation vector per test image using ground truth information—rather than the amortized inference approach employed by our model. This fundamental methodological difference undermines a direct comparison, though we include these results for completeness.
As for generation, our model enables seamless reconstruction at arbitrary resolutions. Figures 9 and 10 show results from LDMI trained on CelebA-HQ at $(256\times 256)$ and $(64\times 64)$, respectively. To achieve this, we encode the observations, sample from the posterior distribution, and transform these latent samples into INRs using our HD decoder. The resulting INRs are then evaluated on grids of varying resolutions to produce reconstructions ranging from ×$\times 1/8$ to ×$\times 4$. Importantly, such flexible resolution control is not supported by methods like GASP, which cannot perform reconstructions due to their adversarial design.
Model Chairs (dB) $\uparrow$ ERA5 (dB) $\uparrow$
Functa (Dupont et al., 2022a) 29.2 34.9
@@ 4.3 Reconstruction
The dual capability of high-quality generation and accurate reconstruction positions LDMI as a versatile foundation for various downstream tasks requiring both generative and reconstructive abilities, a unique advantage over specialized alternatives optimized for only one of these objectives.
@@ 4.4 Hyper-Transforming
On ImageNet and CelebA-HQ at $256\times 256$, our LDMI was trained using the hyper-transforming strategy. As discussed in Section 3.5, instead of training the model from scratch, we reuse the encoder and diffusion model from a pre-trained LDM, training only the Hyper-Transformer Decoder. This approach leverages both the structured latent space learned by the autoencoder and the diffusion model already fitted to it. Training then focuses solely on transforming latent codes into the function space of INR parameters, enabling efficient specialization without retraining the generative backbone.
Our method is agnostic to the latent encoder—VAE, VQVAE, or VQGAN—highlighting the flexibility of the hyper-transforming framework. These results demonstrate its scalability for INR-based synthesis and adaptation of pre-trained LDMs to resolution-agnostic generation.